# On the Limit of Language Models as Planning Formalizers

### v1

@@ 1 Introduction
This paper explores the limit of LLM-as-formalizer devoid of the above two simplifications. We use LLMs to generate the entirety of a PDDL representation, including the domain file and the problem file, given a natural-sounding description of the environment and the task (see Figure 1). On one of the most widely used planning datasets, BlocksWorld PIC (1998) and its derivative, MysteryBlocksWorld Valmeekam et al. (2024), we benchmark a suite of LLMs on generating PDDL that is both solvable and correct. As the descriptions in these datasets are templated, we also construct model-generated, human-validated descriptions that are natural-sounding to different levels.
@@
  • On fully-observed environments such as the BlocksWorld domain, larger GPT models can decently generate entire PDDL, while smaller open-source models cannot.
@@
  • Errors in PDDL generation span syntax and semantics in both the domain and problem.
@@ 2 Task: Formal Planning with PDDL
Formal planning (or classical planning) with PDDL involves a domain file ($\mathbb{DF}$) and problem file ($\mathbb{PF}$) (Figure 1). $\mathbb{DF}$ describes general properties in a planning domain that holds true across problems, while $\mathbb{PF}$ describes specific configurations of each problem instance. Concretely, the $\mathbb{DF}$ defines all available actions for a state-based environment as well as predicates that represent the properties of object types. Each action definition contains the name of the action, parameters and semantics. The semantics of an action include the preconditions which describe the necessary world states where the action is valid to execute, and effects which describe how the states change after the action is executed. The $\mathbb{PF}$ defines the involved objects, the initial states, and the goal states. These two files are then given to a deterministic planner which will algorithmically search for a plan. Such a plan is a series of executable, instantiated actions that sequentially transition the world states from initial to goal. In the AI community, classical planning has been deemed an effective approach to solve real-world users’ problems, as the process is precise, explainable, verifiable, and deterministic.
@@ 3 Methodology: LLM-as-Formalizer
Figure 2: Methodologies for using LLMs in Formal Planning. LLM-as-Planner generates the plan directly, while LLM-as-Formalizer translates the $\mathbb{DD}$ and $\mathbb{PD}$ into PDDL. Previous work such as Liu et al. (2023a) use the LLM to generate partial PDDL, while we generate the entire PDDL (dashed arrow).
@@
4 Evaluation: Datasets, Models, Metrics
@@
4.1 Data
We consider two datasets, BlocksWorld PIC (1998) and MysteryBlocksWorld Valmeekam et al. (2024), one of the most widely used domains in related work. Here, the model must rearrange stacks of blocks on a table from an initial configuration to a goal configuration using a single arm. While each instance has ground-truth $\mathbb{DF}$ and $\mathbb{PF}$, they are invisible to the models and for evaluation only. Notwithstanding existing datasets, we generate the problem configurations ourselves to maximize control over ablation studies (see Section 5). Problem configurations were generated by randomly sampling the number of blocks between 2 and 15 and number of stacks in the initial state and goal between 1 and the number of blocks. Then, the natural language description $\mathbb{DD}$s and $\mathbb{PD}$s were created using 3 methods for 3 different levels of ‘naturalness’.
Heavily Templated. The $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as MysteryBlocksworld Valmeekam et al. (2024). This description is almost a word-by-word translation of PDDL. For example, for the ‘pick-up’ action, the ground-truth PDDL $\mathbb{DF}$ would be the following:
@@ 4.1 Data
While more natural-sounding than the Heavily Templated version, the description still explicitly discusses the preconditions and effects as well as predicates like ‘clear’.
@@
To generate $\mathbb{DD}$, we ask GPT-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$ for BlocksWorld, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique. We next verify the naturalness of the generated text by making there were variations in language throughout all descriptions generated, but was not giving out unnecessary information.
To generate $\mathbb{PD}$, we first algorithmically and randomly permutate BlocksWorld problems, given various total numbers of blocks and stacks. We then provide the model with a symbolic configuration that contains the number blocks, the initial stack configuration and the goal stack configuration. The model then ‘humanize’ the problem by making it sound natural, given a couple of seed exemplars. We manually verify the correctness of the dataset of the non-templated problems by hand by comparing them against the problem configurations. We then verify the naturalness of the $\mathbb{PD}$ by making sure there is variation but no ambiguity in its language.
@@
In total, we construct 111 problems each varying in the number of objects in the unified domain of BlocksWorld. For each of the two Templated descriptions, there is 1 $\mathbb{DD}$ paired with each of 111 $\mathbb{PD}$s. For the Natural description, there are 111 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s. We refer to this dataset as BlocksWorld-111.
Mystery BlocksWorld is a dataset created in Valmeekam et al. (2024) that obfuscates the original BlocksWorld domain by replacing all the names of the types, predicates, actions, and objects with nonsensical words, akin to a wug test (see an example in Appendix A). This dataset as an control group can effectively test whether models create plans via lexical pattern-matching and memorization. The original Mystery BlocksWorld dataset is Heavily Templated. Since the dataset itself is an artificial perturbation, we do not provide more natural descriptions as that would defeat its purpose. We sampled 100 random problems for our experiment, resulting in 1 $\mathbb{DF}$, 1 $\mathbb{DD}$, 100 $\mathbb{PF}$s, and 100 $\mathbb{PD}$s. We refer to this dataset as MysteryBlocksWorld-100.
Data examples can be found in Appendix A.
@@ 4.2 Metrics
We evaluate the predicted plans following Zuo et al. (2024): solvability and correctness. Solvability indicates the percentage of plans that were able to be found based on the predicted $\mathbb{DF}$ and $\mathbb{PF}$, regardless of whether it can correctly work with the actual environment (in our case, described by the gold $\mathbb{DF}$ and $\mathbb{PF}$). In contrast, correctness is a subset that indicates the percentage of actually correct plans. Solvability was determined using the planner dual-bfws-ffparser implemented by Muise (2016) and Correctness was evaluated using VAL333nms.kcl.ac.uk/planning/software/val.html.
@@ 4.3 Models
For both of the LLM-as-planner and LLM-as-formalizer approach, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b-it, gemma-2-27b-it Team et al. (2024), llama-3.1-8B-Instruct, llama-3.1-70B-Instruct Dubey et al. (2024), gpt-3.5-turbo-0125, gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o1-preview-2024-09-12444platform.openai.com/docs/models o1 is only reported in some experiments due to prohibitive cost.. We query these models using KANI Zhu et al. (2023) with default hyper-parameters. The open-source models are run using 4 RTX A6000 GPUs, averaging about 1062 input and output tokens for the LLM-as-formalizer approach in BlocksWorld-111. To emulate real-life application with minimal user interference, we use zero-shot prompts for all naturalness levels across all datasets (see prompts in Appendix B).
@@ 5.1 Can LLMs formalize?
Models Solvability Correctness
Llama-3.1-8B 0/111 0/111
[…]
Llama-3.1-70B 0/111 0/111
[…]
gemma-2-9b-it 3/111 3/111
gemma-2-9$\dagger$ - 10/111
gemma-2-27b-it 0/111 0/111
gemma-2-27$\dagger$ - 12/111
gpt-3.5-turbo 2/111 1/111
gpt-4o-mini 22/111 3/111
gpt-4$\dagger$ - 7/111
gpt-4o 73/111 66/111
gpt-4$\dagger$ - 37/111
o1-preview 101/111 101/111
o1$\dagger$ - 93/111
Table 1: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on the Natural BlocksWorld-111. Similarly trends hold for the Heavily and Moderately Templated versions (results are shown in Appendix C).
First, we seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, instead of partial components in previous work. Table 1 displays the results on our experiment using the most natural sounding descriptions on the BlockWorld.
These results demonstrate that big enough LLMs can decently generate PDDL on fully observed environments. We can see that for BlocksWorld, the performance for gpt-4o surpasses the performance for gpt-4o-mini, which surpasses the performance for gpt-3.5-turbo. As the model gets larger, the number of syntax errors decreases, indicating larger models’ stronger ability of code generation. This conclusion also holds for the other levels of naturalness (Figure 3).
We also observe that open-sourced models with size up to 70B almost cannot generate PDDL. Llama models cannot able to find a single plan across all three datasets while gemma models show poor though non-zero performance across all three datasets. Detailed analysis of errors will be discussed in Section 5.5.
@@ 5.2 Should LLMs formalize?
Between LLM-as-planner and LLM-as-formalizer, which is the preferred methodology for symbolic planning tasks like BlocksWorld? From Table 1, we can see that the best performing gpt-4o is able to generate solvable PDDL 73/111 times, and of those 73 plans, 66 of them are correct. This far surpasses the LLM-as-planner baseline, which only found correct plans 37/111 times. This conclusion holds for the Templated BlocksWorld-111 data (Figure 3). From Figure 3, for MysteryBlocksWorld-100, we can see that LLM-as-formalizer can generate 70/100 correct plans, which far surpassed LLM-as-planner which did not find a single correct plan as the description becomes unorthodox. For gpt-4o-mini, LLM-as-formalizer still outperforms LLM-as-planner for all cases (all naturalness level across two datasets) except Natural BlocksWorld-111.
These results demonstrate that LLMs as formalizers greatly outperforms those as planners in most cases, whenever these LLMs can foramlize PDDL at all. However, these results also show that models that might not be able to formalize (eg. Llama models) but might still be able to plan, though they will have low performance.
@@ 5.3 The more natural, the harder?
Figure 3: Performance across different naturalness level of description on BlocksWorld-111.
In this section, we discuss the question of whether using humanized descriptions makes the problem more difficult. Results from Figure 3 show that as the problem sounds more similar to PDDL, and less natural, the performance of all the models improves. Also, from our results in the Mystery BlocksWorld domain, LLM-as-formalizer performs quite well for a few of the models. However, looking at the domain and problem descriptions, they are very templated and sound like PDDL. This suggests that a more natural-sounding domain and problem description is much more challenging than templated, less natural sounding descriptions. One potential explanation is that pattern matching a template back to PDDL is much easier than having to first parse all the predicates and objects from a passage. Another reason is a more-natural sounding description may leave out implicit common-sense. For example, the Natural BlocksWorld-111 does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved. However, models often fail to invoke this knowledge and will leave out the ‘clear’ predicate, leading to unsolvable PDDL or incorrect plans.
@@ 5.5 What kind of errors?
Natural BlocksWorld-111
@@
Table 2: Different model errors in generating PDDL for all the natural BlocksWorld-111, manually annotated on a 20–example subset.
@@
Models Syntax Error $\mathbb{DF}$ Error $\mathbb{PF}$ Error
gpt-3.5-turbo 16/20 19/20 2/20
@@
Table 3: Different errors of gpt-3.5-turbo, gpt-4o-mini and gpt-4o on Heavily Templated MysteryBlocksWorld-100, manually annotated on a 20–example subset.
Natural BlocksWorld-111
@@
Table 4: Analysis of errors found in $\mathbb{DF}$ for Natural BlocksWorld-111 and Heavily Templated MysteryBlocksWorld-100 out of 20 randomly sampled instances.
In this section, we discuss the kind of errors in PDDL generation. We perform an error analysis on a random 20 sample subset of problems where a plan was not found, or the found plan was not correct. From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and in the $\mathbb{PF}$. Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect action preconditions and effects, incorrect or missing predicates, and missing or incorrect action parameters. The error analysis can be found for the Natural BlocksWorld-111 in Table 2 and Table 4.
As mentioned before, as the models get larger, the amount of syntax errors in the generated PDDL decreases. Smaller models consistently make mistakes in generating the $\mathbb{DF}$. Interestingly, the most common error made for gpt-4o came from the $\mathbb{PF}$, which is intuitively easier to generate than the $\mathbb{DF}$. Common errors in the $\mathbb{PF}$ included incorrect predicates in the initial state and goal state, possibly because the descriptions are natural rather than templated, so that the implicit information in the $\mathbb{PD}$ may be nontrivial to infer. Regarding the $\mathbb{DF}$, the most common error is an incorrect effect in an action. For example, in the ‘unstack’ action, the model does not make the next block ‘clear’ when the top block has been placed in the hand.
For the open-source models, the most common error is syntax errors. For examples, models repeatedly use the keyword ‘preconditions’ instead of ‘precondition’ which might suggest a lack of grasp of the PDDL language. The smallest gemma model also made many mistakes in syntax, though fewer than the Llama models, However, despite the syntax errors, there are still many semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$, which include missing predicates and incorrect effects in the actions. For the Heavily Templated BlocksWorld-111 results, there is a significant gap between the number of plans that were found, and the number of found that were correct. Upon looking at the generated PDDL, we found that the most common error made was swapping the parameters in the preconditions of the ‘stack’ action, leading to incorrect plans.
For the MysteryBlocksWorld-100 dataset (Table 3), we see a similar trend of the syntax errors decreasing as the models get larger. For all three GPT models analyzed, there are barely any errors in the $\mathbb{PF}$ but rather the most common errors came from the $\mathbb{DF}$. Since the Mystery BlocksWorld domain is a result of lexical perturbation, the task of PDDL generation is akin to symbolic information extraction or translation, devoid of much use of commonsense knowledge. As the descriptions for MysteryBlocksWorld-100 were heavily templated, sounding the most similar to PDDL, all the predicates would be listed out in the $\mathbb{PD}$ and the model would just need to match them to PDDL syntax in the $\mathbb{PF}$. While a similar essence, this is more of a challenge for $\mathbb{DF}$ since the clauses of preconditions and effects are more involved. From Table 4, a similar trend between BlocksWorld-111 and MysteryBlocksWorld-100 also suggests that the LLM-as-formalizer methodology is robust to such perturbation.

### v2

@@ 1 Introduction
This paper explores the limit of LLM-as-formalizer devoid of the above two simplifications. We use LLMs to generate the entirety of a PDDL representation, including the domain file and the problem file, given a natural-sounding description of the environment and the task (see Figure 1). On 4 widely used planning simulations from the International Planning Competition, we benchmark a suite of LLMs on generating PDDL that is both solvable and correct. As the descriptions in these datasets are templated, we also construct model-generated, human-validated descriptions that are natural-sounding to different levels.
@@
  • On various planning simulations, closed-source models like GPT can decently generate entire PDDL, while open-source models like Llama up to 405B cannot.
@@
  • Open-source models succumb to syntax errors unlike GPT models, while semantic errors are common for both types of models.
@@ 2 Task: Formal Planning with PDDL
Formal planning (or classical planning) with PDDL involves a domain file ($\mathbb{DF}$) and problem file ($\mathbb{PF}$) (Figure 1). $\mathbb{DF}$ describes general properties in a planning domain that holds true across problems, while $\mathbb{PF}$ describes specific configurations of each problem instance. Concretely, the $\mathbb{DF}$ defines all available actions (and their parameters and pre- and post-conditions) for a state-based environment as well as predicates that represent the properties of object types. The $\mathbb{PF}$ defines the involved objects, the initial states, and the goal states. These two files are then given to a deterministic planner which will algorithmically search for a plan. Such a plan is a series of executable, instantiated actions that sequentially transition the world states from initial to goal. In the AI community, classical planning has been deemed an effective approach to solve real-world users’ problems, as the process is precise, explainable, verifiable, and deterministic.
@@ 3 Methodology: LLM-as-Formalizer
Figure 2: Methodologies for using LLMs in planning. LLM-as-Planner generates the plan directly, while LLM-as-Formalizer translates the $\mathbb{DD}$ and $\mathbb{PD}$ into PDDL. Previous work like Liu et al. (2023a) use the LLM to generate partial PDDL such as $\mathbb{PF}$ only, while we generate the entire PDDL including $\mathbb{PF}$ and $\mathbb{DF}$. Note the $\mathbb{DD}$ and $\mathbb{PD}$ are always provided and the $\mathbb{DF}$ and $\mathbb{PF}$ are always generated by the LLM.
@@
4 Evaluation: Datasets, Metrics, Models
@@
4.1 Datasets
We consider four simulated planning domains, BlocksWorld, Logistics, Barman from the International Planning Competition IPC (1998), and MysteryBlocksWorld Valmeekam et al. (2024).
BlocksWorld, also used in Liu et al. (2023a), is a domain to rearrange stacks of blocks on a table using a robotic arm. There is 1 type of entities, 5 predicates, and 4 actions.
Mystery BlocksWorld obfuscates the original BlocksWorld domain by replacing all the names of the types, predicates, actions, and objects with nonsensical words, akin to a wug test Berko (1958). This dataset as an control group can effectively test whether models create plans via lexical pattern-matching and memorization.
Logistics, also used in Guan et al. (2023), is a domain to transport packages across different locations using both trucks and airplanes. In this domain, there are 6 types of entities, 3 predicates, and 6 actions.
Barman, also used in Zhu et al. (2024), is a domain to to create cocktails from ingredients using different containers and two robotic arms. In this domain, there are 7 types of entities, 13 predicates, and 12 actions.
Each dataset comes with ground-truth PDDL describing domains ($\mathbb{DF}$) and problems ($\mathbb{PF}$). The input to the model is a natural language description of the domain ($\mathbb{DD}$) and the problem ($\mathbb{DD}$). The output of the model is a plan, namely a sequence of instantiated actions defined in $\mathbb{DF}$. For each of these datasets, the natural language description $\mathbb{DD}$s and $\mathbb{PD}$s were created in 3 different levels of naturalness.
Heavily Templated. For BlocksWorld, Logistics and Barman, the heavily templated $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as MysteryBlocksworld Valmeekam et al. (2024). This description is almost a word-by-word translation of PDDL. For example, for the ‘pick-up’ action in BlocksWorld, the ground-truth PDDL $\mathbb{DF}$ would be the following:
@@ 4.1 Datasets
While more natural-sounding than the Heavily Templated version, the description still explicitly discusses the preconditions and effects as well as predicates like ‘clear’. Moderately Templated datasets are available only for BlocksWorld and Logistics due to the size of Barman.
@@
To generate $\mathbb{DD}$, we ask GPT-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique. We next verify the naturalness of the generated text by making there were variations in language throughout all descriptions generated, but was not giving out unnecessary information.
To generate $\mathbb{PD}$, we provide the model with a symbolic configuration that contains the number blocks, the initial stack configuration and the goal stack configuration. The model then ‘humanize’ the problem by making it sound natural, given a couple of seed exemplars. We manually verify the correctness of the dataset of the non-templated problems by hand by comparing them against the problem configurations. We then verify the naturalness of the $\mathbb{PD}$ by making sure there is variation but no ambiguity in its language.
Figure 3: Performance of both usages of LLMs on Heavily Templated BlocksWorld-100, Logistics-100, and MysteryBlocksWorld-100. Detailed results are shown in Appendix C.
@@
In total, we construct 100 problems varying in complexity for all domains. For each of the two Templated descriptions, there is 1 $\mathbb{DD}$ paired with each of 100 $\mathbb{PD}$s. For the Natural description, there are 100 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s. We refer to these dataset as BlocksWorld-100, MysteryBlocksWorld-100, Logistics-100 and Barman-100. Data examples can be found in Appendix A.
@@ 4.2 Metrics
We evaluate the predicted plans following Zuo et al. (2024): solvability and correctness. Solvability only applies to LLM-as-formalizer and indicates the percentage solvable predicted $\mathbb{DF}$ and $\mathbb{PF}$, regardless of whether the resulting plan can be executed bsaed on the the gold $\mathbb{DF}$ and $\mathbb{PF}$. Correctness indicates the percentage of actually correct plans. Solvability was determined using the planner dual-bfws-ffparser implemented by Muise (2016) and Correctness was evaluated using VAL333nms.kcl.ac.uk/planning/software/val.html.
Figure 4: Performance of LLM-as-planner (as-P) and LLM-as-formalzier (as-F) across different naturalness level of description on BlocksWorld-100 and Logistics-100. Detailed results are shown in Appendix C.
@@ 4.3 Models
For both of the LLM-as-planner and LLM-as-formalizer approach, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b|27b-it Team et al. (2024), llama-3.1-8B|70B|405B-Instruct Dubey et al. (2024), DeepSeek-R1-Distill-Llama-8B|70B Guo et al. (2025) [^4] Due to the cost and zero solvability on our easiest dataset, results of Llama-405B and DeepSeek-8B—70B are ommitted., gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o3-mini-2025-01-31555platform.openai.com/docs/models. We query these models using KANI Zhu et al. (2023) with default hyper-parameters. The open-source models are run using 4 RTX A6000 GPUs, averaging about 1062 input and output tokens for the LLM-as-formalizer approach in BlocksWorld-100. To emulate real-life application with minimal user interference, we use zero-shot prompts for all naturalness levels across all datasets (see prompts in Appendix B).
@@ 5.1 Can LLMs formalize?
We seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, instead of partial components in previous work. Figure 3 displays the results on our experiments using the most natural sounding descriptions on Heavily Templated BlocksWorld-100, Logistics-100 and MysteryBlocksWorld-100. Results on the most complex domain Barman-100 is omitted due to close-to-zero performance for all models.
These results demonstrate that GPT-family LLMs can decently generate PDDL, while open-source models even up to 405B parameters struggle. As formalizer, gpt-4o-mini, gpt-4o, o3-mini demonstrate non-trivial and increasing performance on BlocksWorld-100. On the more complex Logistics-100, gpt-4o-mini succumbs to zero performance whereas the other two show decreased performance. The solvability of gpt-4o-mini is often much higher than its correctness, suggesting a good grasp of the PDDL syntax but a lack of semantic understanding. In contrast, the solvability of gpt-4o and o3-mini is often 80% to 100% of their correctness. On the other hand, open-sourced models can rarely generate PDDL, a low-resource language, despite them being reportedly strong at generating high-resource languages like Python Cassano et al. (2022). All Llama models up to 405B cannot generate any solvable PDDL across all three datasets, while gemma models show poor though non-zero performance on BlocksWorld-100 and Logistics-100, and strong performance on MysteryBlocksWorld-100.
@@ 5.2 Should LLMs formalize?
Between LLM-as-planner and LLM-as-formalizer, which is the preferred methodology? Figure 3 shows that on BlocksWorld-100, gpt-4o is able to generate solvable PDDL 64/100 times, and of those 64 plans, 60 of them are correct. This far surpasses the LLM-as-planner baseline, which only found correct plans 33/100 times. This trend holds for Logistics-100 as well as the Moderately Templated and Natural BlocksWorld-100 data (Figure 4). On MysteryBlocksWorld-100, we can see that LLM-as-formalizer can generate 70/100 correct plans, which far surpassed LLM-as-planner which did not find a single correct plan as the description becomes unorthodox. The superiority of LLM-as-formalizer also extends to gpt-4o-mini but not o3-mini, who shows strong performance as a planner. These results demonstrate that LLM-as-formalizer greatly outperforms LLM-as-planner in most cases, whenever these LLMs can foramlize PDDL at all. However, these results also show that models that cannot formalize (e.g., Llama models) can still plan, though with close-to-zero performance.
@@ 5.3 The more natural, the harder?
We now examine whether using humanized descriptions makes the problem more difficult. Results from Figure 4 show that on BlocksWorld-100 as the problem sounds more similar to PDDL and less natural, the performance of all the models improves. Similar results hold for the Logistics domain (see results in Appendix C). This suggests that a more natural-sounding domain and problem description is much more challenging than templated, less natural sounding descriptions. One potential explanation is that pattern matching a template back to PDDL is much easier than having to first parse all the predicates and objects from a passage. Another reason is a more natural sounding description may leave out implicit common-sense. For example, the Natural BlocksWorld-100 does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved. However, models often fail to invoke this knowledge and will leave out the ‘clear’ predicate, leading to unsolvable PDDL or incorrect plans.
@@ 5.5 What kind of errors?
In this section, we discuss the kind of errors in PDDL generation. We perform an error analysis on a random 20 sample subset of problems where a plan was not found, or the found plan was not correct. From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and in the $\mathbb{PF}$. Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect action preconditions and effects, incorrect or missing predicates, and missing or incorrect action parameters. The error analysis for Natural BlocksWorld-100, Logistics-100 and Heavily Templated MysteryBlocksWorld-100 and can be found in Table 1.
For the open-source models, the most common error is syntax errors on BlocksWorld-100 and Logistics-100. For example, models repeatedly use the keyword ‘preconditions’ instead of ‘precondition’ which might suggest a lack of grasp of the PDDL syntax. Gemma models like gemma-2-27b make significantly less syntax errors (3 out of 20 on BlocksWorld-100) than Llama models like Llama-3.1-70B (20 out of 20), despite being smaller. Despite the syntax errors, there are still many semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$, which include missing predicates. As shown in Table 3, there is a significant gap between the number of plans that were found, and the number of found that were correct. We find that the most common error made was swapping the parameters in the preconditions of the ‘stack’ action, leading to incorrect plans. While making much less syntax errors, GPT models frequently suffer from semantic errors. Interestingly, the most common error made for gpt-4o come from the $\mathbb{PF}$, which is intuitively easier to generate than the $\mathbb{DF}$. Common errors in the $\mathbb{PF}$ include incorrect predicates in the initial state and goal state. Common errors in the $\mathbb{DF}$ is incorrect effects in an action. For example, in the ‘unstack’ action, the model does not make the next block ‘clear’ when the top block has been placed in the hand. For MysteryBlocksWorld-100, there are barely any syntax errors or semantic errors in the $\mathbb{PF}$ but rather the most common errors come from the $\mathbb{DF}$. Since this domain is a result of lexical perturbation, formalizing in PDDL is akin to symbolic information extraction and translation, devoid of much use of commonsense knowledge. Due to the heavily templated descriptions, all the predicates would be listed out in the $\mathbb{PD}$ and the model would just need to match them to PDDL syntax in the $\mathbb{PF}$. While a similar essence, this is more of a challenge for $\mathbb{DF}$ since the clauses of preconditions and effects are more involved. From Table 2, a similar trend between BlocksWorld-100 and MysteryBlocksWorld-100 also suggests that the LLM-as-formalizer methodology is robust to such perturbation.
@@
Natural Logistics-100
gemma-2-9b-it 7/20 20/20 15/20
gemma-2-27b-it 8/20 20/20 20/20
Llama-3.1-8B 20/20 20/20 20/20
Llama-3.1-70B 20/20 20/20 10/20
gpt-4o-mini 2/20 20/20 19/20
gpt-4o 5/20 20/20 19/20
@@
Table 1: Error analysis of LLM-as-formalizer on various datasets, manually annotated on a 20–example subset.
@@
Table 2: Analysis of errors found in $\mathbb{DF}$ for Natural BlocksWorld-100 and Heavily Templated MysteryBlocksWorld-100 out of 20 randomly sampled instances.